# Convergence of Unadjusted Langevin in High Dimensions: Delocalization of Bias

### v1

@@ Appendix D Asymptotic Bias for General Observables
In the proof, we use the notation $x=(x_{1},...,x_{d})\in\mathbb{R}^{d}$.
@@
where in the first identity, we used the notation that ${\bm{\alpha}}=(\alpha_{1},...,\alpha_{4})$ is a multi-index with non-negative entries and $|{\bm{\alpha}}|=4$ implies that $\sum_{i}\alpha_{i}=4$. Moreover ${\bm{\alpha}}!=\alpha_{1}!\cdot...\cdot\alpha_{d}!$ and $\xi^{{\bm{\alpha}}}=\xi_{1}^{\alpha_{1}}\cdot...\cdot\xi_{d}^{\alpha_{d}}$. In the second identity, we used the notation $D_{i}^{k}u=\frac{\partial^{k}}{\partial x_{i}^{k}}u$ and the fact that $\mathbb{E}[\xi_{i}^{2}]=1$ and $\mathbb{E}[\xi_{i}^{4}]=3$ for $\xi_{i}\sim\mathcal{N}(0,1)$. In the third identity, we performed the Taylor expansion at $x$, based on the fact $\bar{x}=x+h\nabla\log\pi(x)$. In the last identity, we noticed the fact that the terms in the big bracket equals $\Delta^{2}u$.
@@
$=\frac{1}{2}h\int\pi\Big{(}\nabla f\cdot\nabla\log\pi+\frac{1}{2}|\nabla\log\pi|^{2}_{2}f\Big{)}$
@@
where in the last identity, we used the fact $\pi\nabla\log\pi=\nabla\pi$ and integration by parts.
@@
(D.11) $\left|\int f(\Delta\log\pi)\pi\right|=\left|\int f(\Delta U)\pi\right|\leq\sqrt{\int(f)^{2}\pi}\sqrt{\int(\Delta U)^{2}\pi}\leq\sqrt{C_{0}}\sqrt{\int(f)^{2}\pi}\,.$
When $f$ only depends on $K$ number of coordinates of $x\in\mathbb{R}^{d}$, then the integral $\int(f)^{2}\pi$ will only scale with that number, rather than $d$, because the whole integral will only rely on the marginal distribution of $\pi$ at these coordinates. As such argument also applies to the term $-\frac{1}{4}\int(\Delta f)\pi$, we get that the first order term of the bias depend only on $K$ rather than $d$.

### v2

@@ Appendix D Asymptotic Bias for General Observables
In the proof, we use the notation $x=(x_{1},...,x_{d})\in\mathbb{R}^{d}$. The assumptions used in Proposition 3.1 are
  • Both $u,f$ are smooth, so that the pointwise Taylor expansion is valid.
  • $\lim_{h\to 0}\frac{1}{h}\int\pi(\mathcal{L}_{h}u-\mathcal{L}u)=\int\pi(\lim_{h\to 0}\frac{\mathcal{L}_{h}u-\mathcal{L}u}{h}).$
  • $\lim_{h\to 0}\frac{1}{h}\int(\pi-\pi_{h})(\mathcal{L}_{h}u-\mathcal{L}u)=0.$
These assumptions are made for technical reasons and arise naturally in the Taylor expansion calculations. They may be verified on a case-by-case basis. Here, for simplicity, we assume they hold and use them to derive the asymptotic expansion of the bias of certain observables.
@@
where in the first identity, we used the notation that ${\bm{\alpha}}=(\alpha_{1},...,\alpha_{4})$ is a multi-index with non-negative entries and $|{\bm{\alpha}}|=4$ implies that $\sum_{i}\alpha_{i}=4$. Moreover ${\bm{\alpha}}!=\alpha_{1}!\cdot...\cdot\alpha_{d}!$ and $\xi^{{\bm{\alpha}}}=\xi_{1}^{\alpha_{1}}\cdot...\cdot\xi_{d}^{\alpha_{d}}$. In the second identity, we used the notation $D_{i}^{k}u=\frac{\partial^{k}}{\partial x_{i}^{k}}u$ and the fact that $\mathbb{E}[\xi_{i}^{2}]=1$ and $\mathbb{E}[\xi_{i}^{4}]=3$ for $\xi_{i}\sim\mathcal{N}(0,1)$. In the third identity, we performed the Taylor expansion at $x$, based on the fact $\bar{x}=x+h\nabla\log\pi(x)$. In the last identity, we noticed the fact that the terms in the big bracket equal $\frac{1}{2}\Delta^{2}u$.
@@
Note that in the above, going from the pointwise result (D.3) to the averaged result (D.5) rigorously requires some justification. Here, it is guaranteed by assuming
$\lim_{h\to 0}\frac{1}{h}\int\pi(\mathcal{L}_{h}u-\mathcal{L}u)=\int\pi\lim_{h\to 0}\frac{\mathcal{L}_{h}u-\mathcal{L}u}{h}\,.$
@@
$=\frac{1}{2}h\int\pi\Big{(}\nabla f\cdot\nabla\log\pi+\frac{1}{2}|\nabla\log\pi|^{2}_{2}f\Big{)}+o(h)$
@@
where in the first line, we have used the assumption that $\int(\pi-\pi_{h})(\mathcal{L}_{h}u-\mathcal{L}u)=o(h)$. In the last identity, we used the fact $\pi\nabla\log\pi=\nabla\pi$ and integration by parts.
@@
(D.11) $\left|\int f(\Delta\log\pi)\pi\right|$ $=\left|\int f(\Delta U)\pi\right|$
$=\left|\int f(\Delta U-(\int\Delta U\pi))\pi\right|$
$\leq\sqrt{\int(f)^{2}\pi}\sqrt{\int(\Delta U-(\int\Delta U\pi))^{2}\pi}$
$\leq\sqrt{C_{0}}\sqrt{\int(f)^{2}\pi}\,.$
When $f$ only depends on $K$ number of coordinates of $x\in\mathbb{R}^{d}$, then the integral $\int(f)^{2}\pi$ will only scale with that number, rather than $d$, because the whole integral will only rely on the marginal distribution of $\pi$ at these coordinates. As such argument also applies to the term $\frac{1}{4}\int(\Delta f)\pi$, we get that the first order term of the bias depend only on $K$ rather than $d$.